OpenStudy (anonymous):

What is the y-intercept of the graph of the equation? 2y+6=5x A -6/5 B 3 C 5/2 D -3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Replace x with 0 and solve for y 2y+6 = 5x 2y+6 = 5*0 2y+6 = 0 2y+6-6 = 0-6 ... Subtract 6 from both sides. 2y+0 = -6 2y = -6 2y/2 = -6/2 ... Divide both sides by 2 to isolate y. 1y/1 = -3 1y = -3 y = -3 So the y-intercept is -3. What does this mean? It means that the graph crosses the y axis at the point (0,-3)

OpenStudy (anonymous):

Thanks again, you are a lot of help :)